From: The Southern Quaterly, vol. 39, no. 4, 2001, p. 58-62.

Tragedy on the road to Enigma: Ritual death in Harry Crews’s The Gospel Singer

Anne FOATA Université Marc Bloch, Strasbourg, France

Just as Dionysus is he who devours and is devoured, so Apollo is he who pursues and flees.

Viewed from the panoramic vantage afforded by thirty odd years of novel writing, Harry Crews’s violent, macabre, often farcical world of grotesques is apt to convey an unsettling impression of sameness. Unsettling, that is, for the writer of the present article, a dedicated reader of Crews’s novels, who might appear not to do justice to the enormous variety of characters in his fiction and, hence, to his sustaining ability to create life-like, albeit extreme, people and situations.

But sameness all the same, if one may say so, to the extent that any one of Crews’s characters is prey to an obsession; and who of all people can be more monochord, monobasic, and let’s say it, monotonous than a monomaniac? Karatekas, bodybuilders and beauty queens share the same obsessive worship of perfection, car freaks consume their vehicles in manic communion, compulsive sinners hanker after salvation. Most of Crews’s males lust after female flesh and about all of his characters long to be somewhere else, far from their ordinary lives and ordinary selves, in a frenzied, misdirected search for meaning.

THE GOSPEL SINGER

Thus, Crews’s grotesque Southern landscape may appear as the landscape of modern man’s soul set adrift in a world deprived of the old Christian certainties and prey to untrammeled desires of every kind. This vision of deformed, i.e., depraved, humanity may be the sour milk the young Crews suckled in the stern Calvinistic surroundings of his native Southern Georgia, compounded by the ignorance and superstition of its tenant farmers, the squalid poverty and utter destitution of the Depression years during which he grew up. It is a vision all too familiar in many a Southern novel of its kind, the very stuff that must have boomed down on many a chastened audience during revivals in the Bible Belt. It was, however, for the Roman Catholic Flannery O’Connor, Crews’s fellow Georgia writer, to articulate its grassroots metaphysics most convincingly.

In O’Connor’s words, Crews’s outlook on life is the “prophetic vision” of the “realist of distances” who does not hesitate to distort appearances in order to show a hidden truth to a world that is no longer willing or able to see. He is fond of quoting her when she said: “For the hard of hearing you shout, for the very nearly blind you draw large pictures”(Getting Naked 29), adding in his own words that “if there are freaks in my novels… it is only that these people have conditions which are more apparent and more immediate than the people around them” (ibid 30).

The freak, then, for Crews as he was for O’Connor, is but “the figure of our essential displacement” (Mystery 45) from the basic values of Christianity that have shaped the Western world (so far). Told differently, we are all of us grotesques of some sort in our alienation from God, but as our spiritual deformity has become the condition of the world, we no longer perceive it as such, and think we are normal.

Crews’s “job” (as he says) over the years has been to jolt the reader out of his comfortable belief and make him see “grotesquerie as grotesque and freakishness as freakish” (Getting Naked 29), which, looking back on the thirty years or so of his novel writing, he has certainly done, with a vengeance!

All this came back to me when some time ago I chanced upon an old Time magazine review of Leslie Fiedler’s book on Freaks that I have kept in my Crews files. Together with a picture of the wedding feast from Director Todd Browning’s 1932 film Freaks, it shows a series of six woodcuts reproduced from the late mediaeval Liber Chronicarum of 1494 that probably appeared in Fiedler’s book.

There, staring at me amidst the six “nightmare confrontations with the body” (the picture’s caption), was the image of Foot, the freakish alter ego of Crews’s Gospel singer, with his twenty-seven-inch-long podiatric appendix, “the biggest foot in the world” (The Gospel Singer 29); and flowing into my memory with his (its?) characteristic squirming movement (“like a huge caterpillar”) also came The Thing, his arms looking like legs and his legs like arms and his huge square head jammed below his collarbones, while from far back resounded the screeching of the chicken being torn up and eaten alive by the Geek.

ANNE FOATA

Rereading The Gospel Singer on the urge of these images, I was struck again by the stark inevitability of the story, the black, unremitted vision of life and humanity it exudes, the awesome catharsis of its ending. And full of admiration, nay awe, once again in front of the power, the lean (despite all baroque excess) perfection of Crews’s creation, the spare tension and terseness of his prose. This being all the more so that some of the novels that have succeeded The Gospel Singer over the years had somehow failed to convince me as to the appropriateness of their plot lines and the effectiveness of their closures. Crews’s weakness, it seemed to me - barring a few very good ones- lay in his endings, in his failure to provide convincing, “final” dénouements, which in turn reflected upon the somewhat far-fetched and gratuitous character of some of his novels.

Certainly not so in The Gospel Singer, however. Here was catharsis in the old Aristotelian meaning of the word, here were characters driven by some god-inflicted madness, here was hamartia as flawed and tragic as you may want it, the bleak ineluctability of death and retribution, the strict husbanding of means toward end. Here, in short, was Tragedy.

If one follows the lead of its title and confines oneself to the world of Gospel singing, soul saving and the laying on of hands, one might expect The Gospel Singer to be what it purports to be, a tragedy of the Western world, where the lame and the halt, the freaks of the Freak Fair and the Gospel singer himself are but God’s punishment for the world’s sins and the “signposts for humanity“ (Gospel 44). This seems further enhanced by the epigraph Crews added to the novel, namely: “Men to whom god is dead worship one another.”

There are clues, however, embedded in the narrative that hark back to the rituals of a much older world, to tragedy’s own beginnings, when the orgiastic killing and consumption of vegetation gods had not yet given rise to their apotropaïc, i.e., propitiatory, representation on the tragic stage.

Did Crews have that whole mythic background in mind when he set the action of The Gospel Singer in Enigma, named a character Didymus (“Hell of a name for a man of God to have”, 50), and set about to follow the strict rules of unity of purpose and reversal of fortune enacted by the great Greek tragic playwrights and later codified by Aristotle? Did he mean his unnamed Gospel singer to be murdered and sacrificed like Adonis or Attis of yore so that rain can fall and vegetation grow? He may not have done so consciously, but as newspaperman Richard Hognut (one of the new priests of our age of medias) has it, “there is nothing so predictable as the ritual of catastrophe and tragedy” (198). Both go far back into the ages of mankind and bear the color of the human heart, their recurrence through history merely clad in the vestments of their time.

Thus the Gospel singer’s woman-shy, self-flagellating manager, Didymus (“the violent murderous lover of God,” 58), might be an avatar of one of the Galli, whom the Greeks called Corybantes, those castrated priests of the Phrygian goddess Cybele, who commemorated the slain Attis’ death and resurrection on a wild rampage of mourning and violence to the sound of savage music, while they inflicted on themselves and any unwary onlooker those bloody mutilations no newspaper or network was yet to report on the next day. Or, couldn’t we see in the infuriated crowd that hounds the Gospel singer the avenging Erinyes that pursued Orestes after the murder of his mother? After all, hasn’t the Gospel singer murdered both his childhood sweetheart and his black childhood pal, albeit indirectly?

The unleashing of violence is one of the few constants of mankind. With the orgiastic frenzy of the crowd of Enigma and the lynching of the hero, it seems we plunge again into the murky waters of the worship of Dionysus with its flesh-tearing, flesh-eating followers, and eventually into the ritual death and dismemberment of the god himself in the Orphic version of Dionysus-Zagreus.

At first, however, the general characterization of the Gospel singer, the very name of his manager Didymus would more pointedly outline the far less tragic figure of Apollo, sunny triumphant Apollo, forced by the end of the novel to stumble into the dark revelries of the god who shared his shrine at Delphi.

Like the musician God whom Homer shows playing the lyre to the assembly of the Olympians, the Gospel singer’s radiant good looks and golden hair make him shine like Apollo Phoebus, the sun god, who was also Xanthus, the “fair,” “he of the golden locks”, Apollo Chrysocomes. To the crowds gathering to hear him sing, he is also the healer of bodily afflictions (Apollo Alexikakos), the saviour of diseased souls and parched fields, the rainmaker, as Apollo must have appeared in one of his earler avatars, before later Greeks promoted him to the Pantheon of their gods. By the end of the novel, he has turned into a second remove murderer, not unlike the Hecatebolos who shot his arrows from afar. He also shares the god’s gift for divination and prophecy, harassed as he is by his supplicants who demand that he predict and shape their future, divine their innermost longings and fulfill them.

At this point, the name of Didymus finds its proper niche, not a man’s but a place’s, the very toponym of one of Apollo’s earlier oracles in Asia Minor, before he won the battle against Python and settled in glory at Delphi. Couldn’t we fancy the Gospel singer’s manager as one of the priests of Apollo’s oracle at Didymus, “furiously writing” in his Dream Book?^{1} As an aside and keeping in mind Didymus’ “furious” writerly industry, might not Crews have been elated in discovering the real-life first century (B.C) Alexandrian scholar of this name, whose assiduous writing of reputedly three thousand and five hundred books earned him the nickname of “Chalkenteros” or “Brazen Guts”?

To come back to the Gospel singer, why should his compulsive need for penance after every carnal sin remind one so much of Apollo’s self-exile to the Thessalian Vale of Tempe, where he sought to expiate the killing of the serpent Python: the Serpent, that age-old archetype of promiscuous coupling, the lustful arch-seducer of the Book of Genesis?

For the Gospel singer is weak in the flesh, aroused by every passing tit, rounded hip, and promise of the eye. This in fact is his hamartia, the tragic flaw that leads to his undoing, as it most certainly gets into trouble some of the fifteen-year-old virgins that he seduces underneath the stages of the gospel singing circuit or on the back seat of his Cadillacturned-brothel. Many too were the adventures of fair Apollo, and many the maidens, princesses and nymphs driven to a bitter end: Daphne turned into a laurel, Castalia throwing herself into a fountain, and poor Coronis, the Lapith princess, burned on a pyre by the jealousy of Apollo’s twin, Artemis.

Had the Gospel singer stayed clear of Enigma and its man-devouring Sphynx, Crews’s novel might have read as a paean to gospel singing indeed,^{2} with the protagonist’s progress across the United States no less triumphant than the procession of Apollo and his Muses around Greece. Enigma, however, effaces the golden image of the Musagetes, or, perhaps by way of the sacrificed Adonis, that other golden swain, makes him endorse the fatal lineaments of Dionysus’ story. Enigma, in other words, turns the paean to Apollo into a wild, lethal, dithyramb for Dionysus. Tragedy catches up. Enigma, indeed, as the very first sentence of the novel informs the reader, “is a dead end.”

1 Didymus’ Dream Book is further reminiscent of the Sibylline Books, the books of the Sibyl of Cumae near Naples, which were kept in the temple of the Capitol in Rome.

2 and not, of all things, to sinning, as a blurb on my Dell edition of the novel proclaims, “A powerful paean on the wages of sin” (from the Milwaukee Journal)!

Clad in the long white robe of his trade, his golden hair curling around his god-like face, the very image of the effeminate Dionysus wearing the long Lydian robe, the Gospel singer in Woody Pea’s revival tent strikes the crowd of his worshippers with the raving, destructive madness the god once wreaked upon his followers, in Thebes, Attica, Argos, wherever he went. Like the Bacchants and Bacchantes on the rampage on Mount Cythaeron near Thebes, they bite and tear and claw, the Gospel singer’s mother no less ferocious than King Pentheus’ own. Old men lose their footing, babies thud from their mothers’ laps, crutches, braces and wheelchairs flow in all directions, while the cripples go under the raging mob and Didymus collapses “like an empty sack” (194), clutching his Dream Book.

The collective hysteria ends in a nearby field, with both the Gospel singer and the black preacher hanging from the same tree. Tragedy has been consummated. “The thousands of faces” that had been feeding on his words, on his singing, “but mostly feeding on him… on the beauty of his face and the beauty of his voice” (56-57), have become one huge collective mouth to swallow the god alive.

As in the Orphic version of the myth, and to paraphrase Plutarch, the god has been destroyed, he has disappeared; he has relinquished life and… been born again.

Born again? Yes indeed, as Mirst and Avel, the Gospel singer’s brother and sister, are seen to take over in the last pages of the novel and about to launch into the Gospel singing career with a far-seeing manager of their own. “Le roi est mort, vive le roi!” went the official proclamation at the Court of France at each demise of a king, the person himself far less important than the function he embodied in the light of perennity.

Here the reader who has seen French Director Marcel Camus’s 1959 film Orfeu Negro set in Brazil during the Rio Carnival will perhaps remember the young favella boy hurrying to retrieve the slain Orfeu’s guitar so that he may sing the sun to rise on another day, and life to go on. And we may leave the Gospel singer on this last image of yet another tragic hero, the poet and musician Orpheus rent to pieces by the Thracian women for failing his own MaryBell Carter.

Tragedy comes dressed in many a garb, but tragedy is one in its essence. The Greeks, who had a sunnier notion of man’s carnality, may not have called their gods’ wrath upon the Gospel singer’s overindulgence of the flesh, but they knew about Transgression and made it the essence of Tragedy.

In The Gospel Singer, Harry Crews has reached far beyond the Christian belief in a Redeeming God to remind his readers of the ritual killing of the pharmacos so that the city, or the community, may erect a new safety zone and go on living its ordinary life. Sacrifice, the outcome of what René Girard has called “la violence fondatrice“ (“generative violence“), ensures that the community of men can find an area in which to live without fear of violence and death… for some time at least.

Law and order have been restored to Enigma. Violence has once again faded in the background. But the survivors keep mum about it. And they are so keen on erasing it from their memory, so eager to clean the slate, that in time it will erupt again, somewhere, somehow. As Harry Crews has shown in not a few of the novels that have succeeded The Gospel Singer.

Bibliography

Roberto Calasso. Le Nozze di Cadmo e Armonia. Milano: Adelphi, 1988

(in the present author’s translation from the French.) Harry Crews. The Gospel Singer. New York: William Morrow, 1968.

(Quotations are from the 1969 Dell Paperback, which the novelist has inscribed to the present writer.)

. A Childhood: The Biography of a Place. New York: Harper

& Row, 1978; reissued by U. of Georgia Press, 1995.

Anne Foata. “Interview with Harry Crews”, RANAM (U. of Strasbourg), 5, 1972; 207-225; repr. in Getting Naked with Harry Crews, ed. Erik Bledsoe. The University Press of Florida, 1999; 26-48.

René Girard. Violence and the Sacred. Baltimore: The Johns Hopkins U. Press, 1977).

New Larousse Encyclopedia of Mythology, with an Introduction by Robert Graves and translated from the French by Richard Aldington and Delano Ames. London: Hamlyn, 1959. (Greek mythology, 84-198; more particularly: Apollo, 109-120; Dionysus, 154-163; the Plutarch citation is on 160).

Flannery O’Connor. Mystery and Manners, eds. Sally and Robert Fitzgerald. New York: Farrar, Straus, 1969.

R.Z. Shephard. “Leslie Fiedler’s Monster Party”, Time, 20 Feb. 1978; 95-96. (on Leslie Fiedler’s Freaks. New York: Simon & Schuster, 1978).